# Multitask methods for predicting molecular properties from heterogeneous data

### v1

@@ IV.2.2 Comparison with $\Delta$-learning
variance $1.1\times 10^{-5}$ $3.2\times 10^{-5}$
Table 1: Water Trimer case. Summary statistics for absolute quantum chemistry predictions and for differences.
@@
$\Delta$-learning is trained on differences between predictions of our quantity of interest while the single and multitask methods are trained on absolute predictions of the quantity of interest. Table 1 reports summary statistics for these data sets. In this case, the mean and variance of the differences between quantum chemistry methods are smaller by at least an order of magnitude than those of any set of quantum chemistry predictions. Since $\Delta$-learning is trained on differences to predict differences, it is expected that this method will produce smaller mean absolute error than a method trained on absolutes to predict absolutes.
@@
By modifying our choice of tasks, we can design a multitask model which trains on and predicts differences. We call this the multitask $\Delta$ option. Fig. 5 compares this option to the multitask method on absolutes as well $\Delta$-learning. For each inference method, the average mean absolute error obtained for six random assignments of the training data is plotted for each of the same data set sizes which were considered in our previous experiments. The CAT secondary training set structure is used for the multitask absolute models as well as the multitask difference models. As expected, the multitask absolute model is not as successful as those which train on and learn differences.
@@
In subplot (a) of Fig. 5, the object of estimation is the difference between CCSD(T) and PBE computations for each molecular system. For the multitask $\Delta$ option, we let the primary task be exactly this difference. The secondary task is the difference between SCAN and PBE computations which has correlation coefficient $0.782$ with the primary task data. Results of the multitask $\Delta$ option are plotted in gold. The basic $\Delta$-learning model is plotted in indigo and uses only two levels of theory, CCSD(T) and PBE, to predict their difference. Note that these models are the single task variant of the multitask $\Delta$ approach—they are trained only on the primary task. Therefore, Subplot (a) provides another example where the multitask method outperforms its single task counterpart, in this case with correlation parameter $\rho=0.782$. The turquoise points represent a $\Delta$-learning model trained with three levels of theory: CCSD(T), SCAN, and PBE. This model is the sum of a pair of two level models where the top level of the first difference—between SCAN and PBE—is the bottom level of the second—between CCSD(T) and SCAN. Given sufficiently large training sets for the lower levels of theory, this method is comparable with the multitask $\Delta$ option. The multitask $\Delta$ option outperforms most other models, particularly when either C or A is relatively small.
@@
Subplot (b) of Fig. 5 shows the performance of the same set of methods when the primary and secondary tasks demonstrate a slight negative correlation. In this case, the extra data from the secondary task does not contribute to improved inference of the primary task. Here, we estimate the difference between CCSD(T) and SCAN predictions. For all descriptions accompanying Subplot (a), we reverse the roles of SCAN and PBE for Subplot (b). The major difference between the cases is the correlation coefficient between data for CCSD(T)-SCAN and PBE-SCAN: $\rho=-0.271$. The multitask models shown in Subplot (b) demonstrate the limitations of choosing Pearson’s coefficient as the correlation parameter. The multitask $\Delta$ option and three level $\Delta$-learning do not perform as well as two level $\Delta$-learning in this setting. This behavior suggests that the multitask method requires a stronger correlation between tasks to outperform the single task method. Further, the subplot shows that the multitask $\Delta$ option outperforms three level $\Delta$-learning for models trained on larger amounts of DFT data. Thus, three level $\Delta$-learning seems to suffer more from problematic data than the Multitask $\Delta$ option. Importantly, $\Delta$-learning has no correlation parameter and cannot adjust the contribution of each difference model. Though we have left a detailed optimization of the Multitask $\Delta$ option’s correlation parameter for future work, we observe that $\rho=0$ can always be chosen. Therefore, the fully optimized multitask $\Delta$ option will always perform at least as well as two level $\Delta$-learning. Consequently, the multitask framework makes it possible to benefit from useful secondary datasets which two level $\Delta$-learning cannot accommodate while offering more protection against uncorrelated data than three level $\Delta$ models.
@@
In practice, we may have more than two tasks available to support inference, so we can further contrast the methods discussed in this section by considering how they scale to more tasks. Suppose that alongside PBE and CCSD(T) we want to include $n$ additional DFT functionals $y_{1},\dots,y_{n}$. Then, $\Delta$-learning requires us to put all quantum chemistry methods in some order—with CCSD(T) first and PBE last and the $y_{j}$ in between. The difference between neighboring pairs of methods is then learned. However, hierarchies in DFA accuracy can be difficult to predict and differ between problem settings. This necessary ordering can thus become rather arbitrary. By contrast, the multitask $\Delta$ option would incorporate the $n$ new secondary tasks based on the differences to PBE, i.e. $y_{j}-\text{PBE}$. Each of these secondary tasks can be directly related to the primary task through a correlation parameter—which thus eliminates the need to impose an arbitrary order.
@@ IV.3 Ionization Potential of Small Organic Molecules
Figure 6: Organic Molecules Case. For different numbers of tasks (indicated by color), plots show MAE versus cost. Note that $10^{-0.3}\textrm{eV}$ is approximately half an electronvolt. For each row, a given fraction of all molecules in the C and A sets are used to generate training data for all secondary levels, and the remaining molecules are split among the secondary levels. The fractions are given at the top of each row. Models represented in column (a) were trained without DFT data from the T set while models in column (b) had access to this data.
@@
In Fig. 6, as for previous results, the reported mean absolute error is the average over six tests produced by randomly assigning molecules to the C, A, and T sets. The left column of subplots compares results for multitask models trained on CA data sets to single task models trained only on CCSD(T) while the right column considers multitask models trained on CAT data. Translucent stripes encompass the results for models with different total numbers of tasks. For the two task case, the secondary task data is generated with PBE, and for the three task case the secondary data comes from PBE and PBE0. In all cases, secondary Target data is only generated with PBE. Recall that in our testing framework, when we train with the T data set, we train a custom model for each Target molecule. Specifically, the only T molecule in the training set is the molecule which we will use the model to predict. We remark that in practice it would not be necessary to train a new model for each molecule in the Target set.
@@
As we observed in the water trimer case, the multitask models consistently perform better than single task models of comparable cost. Especially for relatively small training sets, the multitask method shows an improvement in performance when secondary Target data is included in the training set. Fig. 6 also informs us about the utility of including more than one secondary task in a model. Note that for each number of tasks we consider, we implement models with the same combinations of C, A, and T set sizes. In the top row, molecules in the C and A sets are each assigned to exactly one secondary task, regardless of the total number of secondary tasks. Therefore, in this row, for each model implemented with two total tasks, we consider models of the same cost implemented with three tasks and with five tasks. The results show that for sufficiently large training sets, there is a benefit to training with multiple tasks for a given budget. For smaller datasets, two tasks may lead to performance as good or better than a larger number of tasks. In particular, the top subplot in column (b) suggests that two task models perform particularly well when given a relatively inexpensive training set of the CAT structure. Note that, in each model, only one secondary task is given access to T data. Thus, in this setting, the decrease in performance when new tasks are added likely comes from siphoning off part of the computational budget to train tasks which do not have access to T data.
@@
We can also consider whether it is useful to train on more than two predictions for the same molecular configuration. The middle row of Fig. 6 shows cases where each task has access to training data for the same 50% of molecules in the C set as well as 50% of the molecules in the A set while the remaining molecules are distributed one per task, and the final row shows cases where all tasks have access to all data in the C and A sets. From these subplots, we conclude that there is a benefit to training our model on multiple secondary predictions for the same molecule.
@@ V Conclusion
The present work has demonstrated that inference models see a general benefit from leveraging all available data. Furthermore, use of the multitask framework can reduce the cost of generating entirely new training data by an order of magnitude and can also facilitate the opportunistic incorporation of existing datasets of heterogeneous quality into the training set. Our work is the first to explore a multitask GP regression model in a materials science setting. Significantly, the multitask framework is not restricted to the choice of GP inference, and future work can investigate training models with neural networks and other tools for regression.
@@ Acknowledgements.
Appendix A Training Data Cost Model
@@
model training $\text{model training cost}\ =\ \sum_{j}^{m}n_{j}\hat{c}_{j}$
where $n_{j}$ is the number of training systems from the $j^{th}$ level of theory and $\hat{c}_{j}$ is an estimate of prediction cost for that level of theory. We produce one set of estimates for $\hat{c}_{1},\dots,\hat{c}_{m}$ for the water trimer example and a different set of estimates for the small organic molecules example.
For the water trimers example, we randomly selected ten molecular configurations from our data set as the basis for our cost model. Using Psi4, we computed the 3-b interaction energy at the level of SCAN and CCSD(T) for each configuration. We applied a counterpoise correction to ameliorate basis set superposition error. All computations for the water cost model where performed on physically the same Intel Xeon Platinum 8260 node of MIT SuperCloud (https://supercloud.mit.edu/systems-and-software). All computations for a given molecular configuration where performed on the same basis set subasis set sucore. For this case study, we estimate that a CCSD(T) calculation incurs $37$ times the cost of a DFT calculation based on the average prediction time of these methods on our example molecular configurations.
Several configurations in the small organic molecules data set are larger than the water trimer configurations. The median number of electrons in this data set is $40$, the minimum is $14$, and the maximum is $50$. Due to the expense of ionization potential calculations, we base our cost model on energy calculations for ten randomly sampled configurations with $36$ electrons. Note that CCSD(T) scales as $N^{7}$ where $N$ is the number of electrons in a system while DFT scales as $N^{3}$. Therefore, an estimate of the ratio between the cost of CCSD(T) and the cost of DFT based on $36$ electrons will be smaller than the actual ratio for most molecular configurations in our data set. Furthermore, the prediction of ionization potential was performed using ${\Delta}CC$ (delta coupled cluster). Thus, while our reported results (Fig. 6) estimate that CCSD(T) calculations demand $244$ times the expense of DFT, the true ratio between the cost of CCSD(T) and DFT is even larger.
Appendix B SOAP parameters

### v2

@@ IV.2.2 Comparison with $\Delta$-learning
standard deviation $0.0033$ $0.0057$
Table 4: Water Trimer case. Summary statistics for training data.
@@
$\Delta$-learning is trained on differences between predictions of our quantity of interest while the single and multitask methods are trained on absolute predictions of the quantity of interest. Table 4 reports summary statistics for these data sets. In this case, both the mean and standard deviation of the differences between quantum chemistry methods are smaller by at least an order of magnitude than than the corresponding statistics for any set of quantum chemistry predictions. Greater standard deviation in the distribution of target data tends to result in greater standard deviation in the prediction error of a statistical surrogate which in turn leads to larger MAE. Since $\Delta$-learning is trained on differences to predict differences, it is thus expected that this method will produce smaller mean absolute error than a method trained on absolutes to predict absolutes.
@@
By modifying our choice of tasks, we can design a multitask model which trains on and predicts differences. This reinterpretation of tasks will be referred to as the multitask $\Delta$ option. Fig. 5 compares this option to $\Delta$-learning as well as our original implementation of the multitask model, trained on absolute predictions of the QoI. For each inference method, we report the average MAE obtained over six random assignments of the training data to the C, A, and T sets. We consider the same molecule set sizes which were considered in our previous experiments. The CAT secondary training set structure is used for the multitask absolute QoI models as well as the multitask $\Delta$ option models. As expected, the multitask absolute model is not as successful as those which train on and learn differences.
@@
In subplot (a) of Fig. 5, the object of estimation is the difference between CCSD(T) and PBE computations for each molecular system. For the multitask $\Delta$ option, we let the primary task be exactly this difference. The secondary task is the difference between SCAN and PBE computations which has correlation coefficient $0.782$ with the primary task data. Results of the multitask $\Delta$ option are plotted in gold. $\Delta$-learning predictions for CCSD(T)-PBE are plotted in indigo. Note that these $\Delta$-learning models are the single task counterpart to the multitask $\Delta$ approach—they are a conventional GP model trained only on its primary task. Therefore, Subplot (a) provides another example where the multitask method outperforms its single task counterpart, in this case with a smaller correlation parameter than we considered in Subsection IV.2.1: $\rho=0.782$.
The turquoise points represent a hierarchical machine learning (hML) extension to the $\Delta$-learning model, trained with three levels of theory: CCSD(T), SCAN, and PBE. In this method, the difference between CCSD(T) and PBE is represented as
[…]
$\coloneqq$ $\ \Delta(\text{CCSD(T)},\ \text{SCAN})\ +\ \Delta(\text{SCAN},\ \text{PBE}),$
where the submodels $\Delta$(CCSD(T), SCAN) and $\Delta$(SCAN, PBE) are $\Delta$-learning models as defined in (5). Note that while $\Delta$(CCSD(T), SCAN) can only be trained with molecular configurations in $C$, $\Delta$(SCAN, PBE) can be trained with $C\cup A$. Given sufficiently large training sets for SCAN and PBE, this method is comparable with the multitask $\Delta$ option. The multitask $\Delta$ option outperforms most other models, particularly when either $|C|$ or $|A|$ is relatively small.
@@
Subplot (b) of Fig. 5 shows the performance of the same set of methods when the primary and secondary tasks do not have significant correlation. In this case, the extra data from the secondary task does not contribute to improved inference of the primary task. Here, we estimate the difference between CCSD(T) and SCAN predictions. For all descriptions accompanying Subplot (a), we reverse the roles of SCAN and PBE for Subplot (b). The major difference between the cases is the correlation coefficient between data for CCSD(T)-SCAN and PBE-SCAN: $\rho=-0.271$. The multitask models shown in Subplot (b) demonstrate the limitations of choosing Pearson’s coefficient as the correlation parameter. The multitask $\Delta$ option and hML extension to $\Delta$-learning do not perform as well as standard $\Delta$-learning in this setting. Since standard $\Delta$-learning is the single task counterpart to the multitask $\Delta$ option, this behavior suggests that the multitask method requires a stronger correlation between tasks to outperform the single task method. Further, the subplot shows that the multitask $\Delta$ option outperforms hML for models trained on larger amounts of DFT data. Thus, hML seems to suffer more from problematic data than the Multitask $\Delta$ option. Importantly, $\Delta$-learning and its hML extension have no correlation parameters and cannot adjust the contribution of each level of theory. Though we have left a detailed optimization of the Multitask $\Delta$ option’s correlation parameter for future work, we observe that $\rho=0$ can always be chosen. Therefore, the fully optimized multitask $\Delta$ option will always perform at least as well as standard $\Delta$-learning. Consequently, the multitask framework makes it possible to benefit from useful secondary datasets which standard $\Delta$-learning cannot accommodate while offering more protection against uncorrelated data than hML models.
@@
We conclude our comparison between hML and the multitask $\Delta$ option with a remark about scaling when training data is available for more levels of theory. Suppose that alongside PBE and CCSD(T) we want to include $L$ additional DFT functionals $y_{1},\dots,y_{L}$. Then, an hML model which uses all levels of theory would require that these levels be assigned some order—with CCSD(T) first and PBE last and $y_{1},\dots,y_{L}$ in between. The difference between neighboring pairs of methods is then learned. However, hierarchies in DFA accuracy can be difficult to predict and differ between problem settings. This necessary ordering can thus become rather arbitrary. By contrast, the multitask $\Delta$ option would incorporate the $L$ new secondary tasks based on the differences to PBE, i.e., $y_{\ell}-\text{PBE}$. Each of these secondary tasks can be directly related to the primary task through a correlation parameter—which thus eliminates the need to impose an arbitrary order.
@@ IV.3 Ionization Potential of Small Organic Molecules
Figure 6: Organic Molecules Case. For different numbers of tasks (indicated by color), plots show MAE versus cost. Note that $10^{-0.3}\textrm{eV}$ is approximately half an electronvolt. For each row of plots, $k$ determines the fraction of training data which is shared between any pair of secondary sets. Models represented in column (a) were trained without DFT data from the T set while models in column (b) had access to this data.
@@
|$\frac{\bigl{|}\{\mathcal{C}(s_{n})\cap\mathcal{C}(s_{n^{\prime}})\}\bigr{|}}{|C|}\ =\ \frac{\bigl{|}\{\mathcal{A}(s_{n})\cap\mathcal{A}(s_{n^{\prime}})\}\bigr{|}}{|A|}\ =\ k.$
We provide results for $k\in\{0,\frac{1}{2},1\}$. These cases correspond—in order, from the top down—to the three rows of Fig. 6.
@@
In Fig. 6, as for previous results, the reported mean absolute error is the average result of six tests produced by randomly assigning molecules to the C, A, and T sets. The left column of subplots compares results for multitask models trained using the CA structure to the single task reference while the right column considers multitask models trained with the CAT structure. Translucent stripes encompass the results for models with a given number of tasks, $N$. For the two task case, the secondary task data is generated with PBE, and for the three task case the secondary data comes from PBE and PBE0. In all cases, secondary Target data is only generated with PBE. Recall from Subsection IV.1.4 that our training and testing procedure is designed to produce fair comparisons between models trained on CA and CAT training structures. Though we test the predictive performance of each training set structure on $320$ different target molecular configurations, we do not simultaneously include all of these configurations in $\mathcal{T}(\text{PBE})$. Rather, we iterate over the target configurations and train a multitask model for each configuration such that $\mathcal{T}(\text{PBE})=1$ in each case. We remark that in practice it would not be necessary to train a new model for each molecule in the Target set.
As we observed in the water trimer case, the multitask models consistently perform better than single task models of comparable cost. Especially for relatively small training sets, the multitask method shows an improvement in performance when secondary Target data is included in the training set. Further, we report that for larger secondary training sets, correlation between multitask predictions and the CCSD(T) target data exceeds $95\%$, but we defer a full evaluation of prediction quality at the level of individual molecular configurations to future work.
@@
Fig. 6 also informs us about the utility of including more than one secondary task in a model. Note that we consider the same combination of sizes for the C, A, and T sets regardless of the total number of tasks, $N$. In the top row, $k=0$ implies that molecular configurations in the C and A sets are each assigned to exactly one secondary task, regardless of the total number of secondary tasks. Therefore, in this row, we can compare models implemented with different $N$ but nearly identical training budgets. The results show that for sufficiently large training sets, given a fixed budget, there is a benefit to dividing the budget across multiple secondary tasks rather than employing only one secondary task. For smaller datasets, choosing $N=2$ may lead to performance as good or better than a larger number of tasks. In particular, the top subplot in column (b) suggests that models where $N=2$ perform particularly well when given a relatively inexpensive training set of the CAT structure. Note that, in each model, only one of the secondary tasks is trained using T data: only $\mathcal{T}(\text{PBE})$ is nonempty. Thus, in this setting, the decrease in performance when new tasks are added likely comes from siphoning off part of the computational budget to train tasks which do not have access to T data.
@@
We can also consider whether it is useful to include the same molecular configuration in the training set of multiple secondary tasks. In the final row of Fig. 6, $k=1$ implies that $\mathcal{C}(s_{n})\cup\mathcal{A}(s_{n})$ is the same set for all $n=1,\dots,N-1$. By contrast, in the middle row, to enforce $k=\frac{1}{2}$, we assign $50\%$ of the configurations in $C$ and $A$ to every secondary task while the remaining configurations in these sets are distributed between the secondary tasks so that each is assigned to exactly one secondary task. The improved performance of the $N=3$ and $N=5$ cases in the last row compared with the middle row suggest that there is a benefit to training our model on multiple secondary predictions for the same molecule.
@@ V Conclusion
The present work has demonstrated that inference models see a general benefit from leveraging all available data. Furthermore, use of the multitask framework can reduce the cost of generating entirely new training data by an order of magnitude and can also facilitate the opportunistic incorporation of existing datasets of heterogeneous quality into the training set. To our knowledge, our work is the first to explore an asymmetric multitask GP regression model in a materials science setting. Significantly, the multitask framework is not restricted to the choice of GP inference. Future work can preserve the same relationship between primary and secondary task submodels and investigate training each submodel with neural networks or other tools for regression.
@@ Acknowledgements.
Appendix A Lengthscale and Variance Optimization
Refer to caption Refer to caption
(a) Primary hyperparameters (b) Secondary hyperparameters
Figure 7: Water Trimer Case. Each point compares the test MAE of a GP model which uses the MLE variance to the test MAE of a model which uses a variance estimate obtained by minimizing MAE on the optimization set. Different data points are obtained by shuffling the optimization data, and color indicates the size of the optimization data set. In plot (a) we consider a GP model for the primary regression task while in plot (b) we consider a model $\delta_{s_{n}}$, as defined in (12).
This section provides further detail on our procedure for hyperparameter optimization with a focus on computational tractability. We note that further improvements are possible and worth investigating in future work. Specifically, we discuss optimization for the test case explored in Subsections IV.2.1 and IV.2.2 to estimate the hyperparameters of the prior distributions for $f_{p}$ and $\delta_{s_{1}}$, given by (13) and (14). Both priors are constructed with the squared exponential kernel:
$k(\bm{X},\bm{X}^{\prime})\ =\ v\ \exp\left(\frac{-(\bm{X}-\bm{X}^{\prime})^{T}(\bm{X}-\bm{X}^{\prime})}{2\ell}\right)$
As discussed in IV.1.3, we optimize hyperparameters using a data set of $M=150$ molecular configurations. For all statistical models used in Subsections IV.2.1 and IV.2.2, the lengthscale, $\ell$, is obtained through a maximum likelihood procedure while the variance, $v$, is taken to be the value which minimizes MAE on the optimization set. The latter choice was made because, for the three body energy data, variance estimates produced by minimizing MAE achieve lower test error than the maximum likelihood estimate of variance. Fig. 7 establishes that these choices of $M$ and variance estimate are not necessary for the multitask model to perform well; in fact, the model appears relatively robust to reasonable optimization procedures.
Each color in Fig. 7 represents a different optimization set size, $M$. We obtain multiple hyperparameter estimates for each $M$ by shuffling the optimization data set. For each shuffled realization of the data, we solve for estimates ^$\hat{\ell}_{MLE}$, $\hat{v}_{MLE}$, and $\hat{v}_{MAE}$. The plotted points compare MAE on a test data set of $320$ molecular configurations obtained using ^$\hat{\ell}_{MLE}$ and $\hat{v}_{MLE}$ to MAE for the same test data set obtained using ^$\hat{\ell}_{MLE}$ and $\hat{v}_{MAE}$. We can see that both approaches to estimating variance lead to good predictive performance, but use of $\hat{v}_{MAE}$ generally produces lower MAE than use of $\hat{v}_{MLE}$ for both $f_{p}$ and $\delta_{s_{1}}$. Fig. 7 also demonstrates that the optimization data set size can be reduced without significant reductions in prediction accuracy though there is greater variance in prediction quality for $M\leq 45$.
Note that given ^$\hat{\ell}_{MLE}$, the expression for $\hat{v}_{MLE}$ is closed form and $\hat{v}_{MAE}$ is found using ^$\hat{\ell}_{MLE}$. Thus, we only optimize for one hyperparameter at a time. If the optimization procedure reaches a threshold of 30 seconds, we halt the procedure, reshuffle the data, and restart. In the case that $M=150$, we require on the order of $10$ seconds on a laptop to successfully obtain MLE estimates. MAE estimates are found in $\sim 100$ seconds for the primary task and $\sim 40$ seconds for the secondary task. The reported amounts include any time spent on false starts which time out at $30s$. By comparison, a single DFT calculation for 3-b energy using a SCAN functional approximation requires $\sim 650$ seconds on a cluster. Thus, compared with data generation, hyperparameter optimization accounts for a small portion of the computational budget of our numerical experiments. Our results in Fig. 7 suggest that we can construct successful statistical models even when the budget for generating the optimization data set is much smaller than in our work. For any $M$, the cost of generating the optimization set will be dominated by CCSD(T) computations which are necessary for all statistical models which we implement. Future work may fully evaluate the minimum cost required to produce hyperparameters which will achieve a specified threshold of accuracy and investigate strategies for constructing opportunistic optimization sets from existing data.
Appendix B Training Data Cost Model
@@
model training $\text{model training cost}\ =\ \sum_{\ell}^{L}n_{\ell}\hat{c}_{\ell}$
where $n_{\ell}$ is the number of training systems from the $\ell^{th}$ level of theory and $\hat{c}_{\ell}$ is an estimate of prediction cost for that level of theory. We produce one set of estimates for $\hat{c}_{1},\dots,\hat{c}_{L}$ for the water trimer example and a different set of estimates for the small organic molecules example.
For the water trimers example, we randomly selected ten molecular configurations from our data set as the basis for our cost model. Using Psi4, we computed the 3-b interaction energy at the level of SCAN and CCSD(T) for each configuration. We applied a counterpoise correction to ameliorate basis set superposition error. All computations for the water cost model where performed on physically the same Intel Xeon Platinum 8260 node of MIT SuperCloud (https://supercloud.mit.edu/systems-and-software). All computations for a given molecular configuration where performed on the same core. For this case study, we estimate that a CCSD(T) calculation incurs $37$ times the cost of a DFT calculation based on the average prediction time of these methods on our example molecular configurations.
Several systems in the small organic molecules data set contain more atoms than the water trimer configurations. The median number of electrons per system in the organic molecules set is $40$, the minimum is $14$, and the maximum is $50$. We estimate the cost of level of theory $\ell$ as the computational time required to obtain the total energy at this level of theory averaged over ten randomly sampled configurations with $36$ electrons. Note that this estimate is conservative: CCSD(T) scales as $N_{e}^{7}$ where $N_{e}$ is the number of electrons in a system while DFT scales as $N_{e}^{3}$. Therefore, an estimate of the ratio between the cost of CCSD(T) and the cost of DFT based on $36$ electrons will be smaller than the actual ratio for most molecular configurations in our data set. Furthermore, our prediction of the ionization potential was performed using ${\Delta}CC$ (delta coupled cluster), which requires multiple CCSD(T) calculations to reach an IP prediction. Thus, while our reported results (Fig. 6) estimate that CCSD(T) calculations demand $244$ times the expense of DFT, the true ratio between the cost of CCSD(T) and DFT is even larger. Consequently the reported gains in efficiency for the multitask method are expected to be a lower bound on the true improvement.
Appendix C SOAP parameters